X is 5 years older than y. Z is 6 years older than x. How old are each person if the sum of their aget is 31?
Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20060)

X is 5 years older than Y.
X = Y+5
Z is 6 years older than X.
Z = X+6
...the sum of their ages is 31...
X+Y+Z = 31

First substitute (Y+5) for X in Z = X+6
and simplify.

Then substitute (Y+5) for X and Z = (what you got)
into X+Y+Z = 31 and solve for Y.

Then substitute that value for Y in X = Y+5 to
find X.

Then substitute what you got for X in Z = X+6
to find Z.

I won't tell you the answers but one of them is
twice as old as another of them.

Edwin


Answer by ikleyn(52802)

.
X is 5 years older than Y. Z is 6 years older than X. How old are each person if the sum of their ages is 31?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let "y" is the age of the person Y.

Then the age of X is (y+5), and the age of Z is (y+5)+6, according to the condition.

The sum of ages is y + (y+5) + ((y+5)+6) = 3y + 16, and it is equal to 31.

So you have this equation

3y + 16 = 31.

Solve it for "y" and get the answer.

Should I explain more or you just can to complete it yourself?